OpenStudy (flexastexas):

3. A 12 kg barrel is pulled up by a rope. The barrel accelerates at 1.2 m/s2. Find the force exerted by the rope. Show all your work.

OpenStudy (flexastexas):

I am almost certain I should be using mass * acceleration. I have seen other people solve the problem and they got 12*11= 132 Where did they get 11 from>

OpenStudy (flexastexas):

?

OpenStudy (korosh23):

Did it ask you for which component ? Like x or y?

OpenStudy (korosh23):

maybe 11 m/s^2 is the acceleration of rope that pulls the barrel. Is it given to you?

OpenStudy (flexastexas):

nope

OpenStudy (flexastexas):

But it would be Y because the barrel is being pulled up

OpenStudy (snowsurf):

Since the problem says pulled up by rope this has to imply vertical motion. We need to find the T tension force. Let's draw a free body diagram.  Fg is the force due to gravity and T is the tension force from the rope. Using Newton second law of motion we have \[T - F_g = ma\] \[T - mg = ma\] Fg = mg Solve for T \[T = m (a +g)\] T = (12Kg)(1.2 + 9.8) m/s^2 T = 132 N